# DINO in the Room: Leveraging 2D Foundation Models for 3D Segmentation

### v1

@@ Abstract
Vision foundation models (VFMs) trained on large-scale image datasets provide high-quality features that have significantly advanced 2D visual recognition. However, their potential in 3D vision remains largely untapped, despite the common availability of 2D images alongside 3D point cloud datasets. While significant research has been dedicated to 2D–3D fusion, recent state-of-the-art 3D methods predominantly focus on 3D data, leaving the integration of VFMs into 3D models underexplored. In this work, we challenge this trend by introducing DITR, a simple yet effective approach that extracts 2D foundation model features, projects them to 3D, and finally injects them into a 3D point cloud segmentation model. DITR achieves state-of-the-art results on both indoor and outdoor 3D semantic segmentation benchmarks. To enable the use of VFMs even when images are unavailable during inference, we further propose to distill 2D foundation models into a 3D backbone as a pretraining task. By initializing the 3D backbone with knowledge distilled from 2D VFMs, we create a strong basis for downstream 3D segmentation tasks, ultimately boosting performance across various datasets. Code will be made available upon acceptance.
@@ 1 Introduction
We need to talk about the elephant DINO in the room.
@@
Current state-of-the-art 3D semantic segmentation models use specialized 3D backbones, which are trained from scratch [56, 58, 7]. In contrast, for the highly related task of image segmentation, the current paradigm is to use 2D backbones initialized with pretrained weights of strong vision foundation models  [5, 33, 11, 41, 64, 23, 43, 22] (VFMs). These VFMs are predominantly trained in a self-supervised manner on large-scale image datasets, enabling strong generalization capabilities. However, current 3D datasets [8, 44, 1, 67, 3, 4, 50] are orders of magnitude smaller than their 2D counterparts [46, 47]. As a result, generalist 3D foundation models have yet to emerge. Interestingly, though, we observe that 3D point cloud data is often accompanied by corresponding 2D images [8, 44, 4, 50]. This raises the question: can we leverage the power of 2D foundation models for 3D segmentation?
@@
Figure 1: DINO in the room (DITR). We present a simple approach to (a) inject or (b) distill DINOv2 features into 3D semantic segmentation models that yields state-of-the-art results across indoor and outdoor 3D benchmarks.
@@
For most 3D point clouds, corresponding images are typically available due to the process with which this data is captured. In indoor scenarios, a scene is captured with RGB or RGB-D cameras as a video sequence and a colored point cloud is formed via 3D reconstruction [8, 44]. As a result, the data inherently includes 2D images, and correspondences between images and point clouds are obtained as part of data preprocessing. In outdoor scenarios, 3D street scenes are typically captured with LiDAR scanners mounted on vehicles that are often also equipped with cameras, providing corresponding 2D images [4, 50, 3]. Despite the availability of these images, current state-of-the-art methods for both indoor and outdoor 3D semantic segmentation predominantly focus on 3D data [58, 63, 59, 24]. While 2D–3D fusion has been investigated with moderate success [2, 20, 65, 42, 51], existing works have thus far not capitalized on the existence of semantically rich VFMs.
@@
In this paper, we challenge the current paradigm and claim that VFMs can, as they did in 2D, enrich 3D models with generalist features that are not easily learned from the 3D data. We verify this by showing that a simple combination of a VFM and a state-of-the-art 3D segmentation model [58] significantly outperforms existing models, achieving new state-of-the-art results on several benchmarks. Concretely, we take a frozen DINOv2 ViT [33] and extract image features that correspond to the points in the associated point cloud. Next, we inject these 2D features into their corresponding points in the decoder of the 3D backbone. Then, we train this model for the regular segmentation objective, resulting in the 2D-to-3D injection approach (see Fig. 1 (a)). Notably, the segmentation performance improves significantly, while only using unlabeled images. Especially for indoor datasets with many semantic classes and for outdoor datasets that provide $360^{\circ}$ camera coverage, this setup outperforms the 3D-only baseline on public leader boards by large margins, i.e., $\texttt{+}7.1$ mIoU on ScanNet200 [44] and $\texttt{+}2.4$ mIoU on nuScenes [4]. We call this approach DINO In The Room (DITR).
@@
Moreover, we show that even if images are not available during inference, it is still possible to utilize features from the same frozen DINOv2 model to enhance segmentation performance through distillation [15]. Specifically, we take 3D point cloud datasets that have corresponding images and teach a 3D student model to output features aligned with those extracted from a DINOv2 teacher model, using a distillation objective (see Fig. 1 (b)). Subsequently, the distilled student model can simply be fine-tuned for 3D segmentation in a regular fashion, without requiring corresponding images, and thus without causing any overhead during inference. This distillation pretraining enables the 3D model to capture the semantic richness of 2D foundation models without requiring any labeled data, effectively using them as distillation targets instead of semantic labels. Furthermore, it allows pretraining across multiple datasets without adjusting for dataset-specific semantic label sets, as the target feature space remains consistent. With this setup, we observe consistent improvements on all datasets compared to random initialization. On SemanticKITTI [3], where there is only a single image per 3D LiDAR scan, and on S3DIS [1], where camera views are sparser, this distillation setup even outperforms DITR. We call this approach Distill DITR (D-DITR).
@@
Altogether, these results highlight the significant yet underexplored potential of VFMs for 3D tasks. We hope that our findings will promote the use of VFMs in 3D scene understanding and encourage further research into this direction. Additionally, we note that while we focus on DINOv2 in this paper, as it represents the current strongest VFM, we anticipate that the general setup can leverage even stronger models that we expect to see in the future. Therefore, we recommend that whenever corresponding images are available, they should be used to complement 3D segmentation models, preferably using injection and otherwise using distillation.
@@
  • We demonstrate that DINOv2 can also act as a teacher for pretraining backbones of 3D models through distillation, surpassing the performance of existing unsupervised pretraining methods.
@@ 2 Related Work
3D Semantic Segmentation.  Despite architectural differences, most 3D backbones for dense tasks such as segmentation follow a U-Net-like hierarchical encoder–decoder design with skip connections. Early methods apply point-wise fully-connected layers [38, 39], or continuous 3D convolutions [35, 53] to capture local geometric patterns. Later, MinkUNet [7] significantly improves efficiency and accuracy by voxelizing point clouds and utilizing sparse 3D convolutions [13]. Following the success of Vision Transformers [10], attention mechanisms have been increasingly adopted for 3D segmentation, although their quadratic computational complexity poses challenges for large-scale scenes. To address this, several methods [26, 55, 62, 66, 56, 58] restrict attention to local neighborhoods, where points are subsampled and aggregated hierarchically across multiple layers. Most recently, Point Transformer V3 (PTv3) [58] employs space-filling curves to map 3D point clouds into 1D sequences, effectively preserving the local neighborhood structure while significantly simplifying and accelerating local attention computations. PTv3 achieves state-of-the-art performance across both indoor and outdoor datasets. In this work, we enhance PTv3 with VFM features to further improve performance.
@@
While these works effectively use 2D information for 3D segmentation, we observe that the community has not yet identified or leveraged the power of VFMs. In this work, we demonstrate that a simple injection of features from a strong 2D model—DINOv2 [33]—into a 3D model allows for a significantly better segmentation performance, obtaining new state-of-the-art results. Moreover, DITR is the first fusion approach shown to be effective for both indoor and outdoor scenes, highlighting its general applicability.
2D–3D Distillation.  Recent methods have explored various strategies to distill 2D representations into 3D backbones for enhanced representation learning in LiDAR point clouds. Yan et al. [60] propose a joint training framework, where a 3D model is trained for both 3D segmentation and alignment with a 2D model, improving learned 3D representations. Several approaches apply contrastive learning either on super-points and super-pixels generated by pretrained 2D backbones [45, 29], or on prototypes created in both the 2D and 3D domains [6]. Others [31, 32] propose new contrastive loss functions to address the issue of contrasting semantically similar regions, and conduct extensive experiments with linear probing or limited-data setups. Recently, ScaLR [37] studies the effects of the size of the 3D backbone, 2D backbone, and the dataset in the image-to-LiDAR distillation setup, and shows that scaling the models and pretraining on diverse datasets is crucial for distillation. Interestingly, while these existing methods have leveraged image-to-LiDAR distillation for zero-shot 3D segmentation and linear probing with success, they have not demonstrated that VFM distillation allows for more accurate fine-tuning. In this work, we demonstrate that a simple distillation setup using a strong VFM enables accurate image-free inference with DITR. Moreover, while related works focus on LiDAR data, our D-DITR distillation setup is not only effective in combination with LiDAR data, but also for indoor scenes.
@@ 3 Method
Figure 2: DITR architecture overview. We extract 2D image features from a frozen DINOv2 [33] model  and unproject them (2D-to-3D) onto the 3D point cloud. The unprojected features are subsequently max-pooled to create a multi-scale feature hierarchy. The raw point cloud is fed through a 3D backbone  and the unprojected image features are added to the skip connection between the encoder $\mathcal{E}_{l}$ and decoder $\mathcal{D}_{l}$ block on each level. The model is then trained with the regular segmentation loss.
@@ 3.1 Injection
In many modern 3D datasets, images from calibrated cameras are provided alongside the point clouds [8, 1, 4, 3, 50]. We leverage these images to inject semantically rich 2D features into the 3D backbone. An overview of this process is shown in Fig. 2: we first map 3D points to their corresponding pixels to extract associated DINOv2 features (2D-to-3D Mapping), then inject these features into the skip connections of the 3D backbone’s decoder (3D Feature Fusion).
2D-to-3D Mapping.  Let $\mathcal{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ be a 3D point cloud of $N$ points, and assume a collection of $K$ calibrated cameras, $\{(\mathbf{I}_{k},\mathbf{K}_{k},\mathbf{T}_{k})\}^{K}_{k=1}$. Each camera has intrinsics $\mathbf{K}_{k}$ and world-to-camera extrinsics $\mathbf{T}_{k}$, with $\mathbf{I}_{k}$ denoting its captured image of resolution $H\times W$. We feed each 2D image $\mathbf{I}_{k}$ into a frozen DINOv2 ViT [10], obtaining patch-level embeddings $\mathbf{F}_{k}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D_{\text{2D}}}$, where $P$ is the patch size and $D_{\text{2D}}$ is the feature dimension. For each point $\mathbf{p}_{i}$, we transform it into the $k$-th camera space as $\mathbf{q}_{i}^{k}=\mathbf{T}_{k}\,(\mathbf{p}_{i},1)$ and multiply by the intrinsics $\mathbf{K}_{k}$ to obtain homogeneous pixel coordinates $\left(x_{i}^{k},\,y_{i}^{k},\,z_{i}^{k}\right)=\mathbf{K}_{k}\mathbf{q}_{i}^{k}$. The pixel coordinates are then given by $\left(u_{i}^{k},\,v_{i}^{k}\right)=\left(x_{i}^{k}/z_{i}^{k},\,y_{i}^{k}/z_{i}^{k}\right)$. We consider $\mathbf{p}_{i}$ visible in the $k$-th image if $u_{i}^{k}\in[0,W)$, $v_{i}^{k}\in[0,H)$ and the depth $z_{i}^{k}$ is positive (view-frustum culling). If $\mathbf{p}_{i}$ is visible in the $k$-th image, we determine the corresponding patch index as
@@
3D Feature Fusion.  Because point clouds are unstructured, common 3D backbones [58, 7, 68] voxelize the input, keeping only one point per voxel, thereby reducing $\mathcal{P}$ to $\mathcal{P}^{\prime}$ with $M\leq N$ points. These backbones typically follow a U-Net-like encoder–decoder architecture, processing $\mathcal{P}^{\prime}$ at multiple spatial resolution levels $l\in\{1,\dots,L\}$. At each encoder level, a pooling layer downsamples the point features to a coarser spatial resolution. Formally, let $\mathbf{X}_{l}^{\mathcal{E}}\in\mathbb{R}^{M_{l}\times D_{l}}$ denote the output of the $l$-th level encoder block $\mathcal{E}_{l}$, where $l=1$ is the finest (original) resolution with $M_{1}=M$. After the encoder and the bottleneck $\mathcal{B}$, the decoder upsamples and refines these features back toward the original resolution, yielding $\mathbf{X}_{l}^{\mathcal{D}}$ as the output of each decoder block $\mathcal{D}_{l}$. To simplify notation, we define $\mathbf{X}_{L+1}^{\mathcal{D}}$ to be the output of the bottleneck $\mathcal{B}$.
@@
where ++ denotes element-wise addition, and each $f_{l}^{\cdot}$ is a linear projection into the decoder’s feature dimension followed by batch normalization [19] and GELU [14]. Finally, the output $\mathbf{X}^{\mathcal{D}}_{1}$ of the last decoder block yields per-point features for $\mathcal{P}^{\prime}$, which are passed through a linear segmentation head to produce class logits.

### v2

@@ Abstract
Vision foundation models (VFMs) trained on large-scale image datasets provide high-quality features that have significantly advanced 2D visual recognition. However, their potential in 3D scene segmentation remains largely untapped, despite the common availability of 2D images alongside 3D point cloud datasets. While significant research has been dedicated to 2D–3D fusion, recent state-of-the-art 3D methods predominantly focus on 3D data, leaving the integration of VFMs into 3D models underexplored. In this work, we challenge this trend by introducing DITR, a generally applicable approach that extracts 2D foundation model features, projects them to 3D, and finally injects them into a 3D point cloud segmentation model. DITR achieves state-of-the-art results on both indoor and outdoor 3D semantic segmentation benchmarks. To enable the use of VFMs even when images are unavailable during inference, we additionally propose to pretrain 3D models by distilling 2D foundation models. By initializing the 3D backbone with knowledge distilled from 2D VFMs, we create a strong basis for downstream 3D segmentation tasks, ultimately boosting performance across various datasets.
@@ 1 Introduction
The idiom “elephant in the room” refers to a situation where something important is being ignored, while it should be discussed. Currently, 2D foundation models, such as DINOv2 [36], remain largely ignored for 3D segmentation, despite their ability to provide strong semantic priors. Therefore, we believe that we need to talk about the elephant DINO in the room.
@@
The current paradigm for state-of-the-art 3D segmentation, both in literature and on benchmarks, is to use models with specialized 3D backbones that are trained from scratch [63, 62, 9]. In contrast, for the highly related task of image segmentation, the current paradigm is to use 2D backbones initialized with pretrained weights of strong vision foundation models (VFMs) [7, 36, 13, 44, 71, 26, 47, 25, 24]. These VFMs are predominantly trained in a self-supervised manner on large-scale image datasets, enabling strong generalization capabilities. However, current 3D datasets [10, 48, 1, 74, 5, 6, 55] are orders of magnitude smaller than their 2D counterparts [51, 50]. As a result, while progress has been made [61], generalist 3D foundation models have yet to emerge. Interestingly, though, we observe that 3D point cloud data is often accompanied by corresponding 2D images [10, 48, 6, 55]. This raises the question: how can we leverage the power of 2D foundation models for 3D segmentation?
@@
Figure 1: DINO in the room (DITR). We present an approach to (a) inject or (b) distill DINOv2 features into 3D semantic segmentation models that yields state-of-the-art results across indoor and outdoor 3D benchmarks.
@@
For most 3D point clouds, corresponding images are typically available due to the process with which this data is captured. In indoor scenarios, a scene is captured with RGB or RGB-D cameras as a video sequence and a colored point cloud is formed via 3D reconstruction [10, 48]. As a result, the data inherently includes 2D images, and correspondences between images and point clouds are obtained as part of data preprocessing. In outdoor scenarios, 3D street scenes are typically captured with LiDAR scanners mounted on vehicles that are often also equipped with cameras, providing corresponding 2D images [6, 55, 5]. Despite the availability of these images, current state-of-the-art methods for both indoor and outdoor 3D semantic segmentation typically only use 3D data [62, 70, 64, 27]. While 2D–3D fusion has been investigated with moderate success [2, 22, 72, 46, 56], existing works have thus far not capitalized on the existence of semantically rich VFMs.
@@
In this paper, we challenge the current paradigm and claim that VFMs can, as they did in 2D, enrich 3D models with generalist features that are not easily learned from the 3D data. To verify this, we take a state-of-the-art 3D segmentation model [62] and augment it with 2D VFM features. Concretely, we take a frozen DINOv2 [36] and extract 2D image features that correspond to the points in the associated 3D point cloud. Next, for the 3D points that have been matched to 2D features, we inject these 2D features into the 3D model at different decoder stages. Then, we train this model for the regular segmentation objective, resulting in a 2D-to-3D injection approach (see Fig. 1 (a)). With this injection approach, which only uses unlabeled images, the semantic segmentation performance improves significantly, achieving new state-of-the-art results. Especially for indoor datasets with many semantic classes and for outdoor datasets that provide $360^{\circ}$ camera coverage, this setup outperforms the 3D-only baseline on public leaderboards by large margins, i.e., $\texttt{+}7.1$ mIoU on ScanNet200 [48] and $\texttt{+}2.4$ mIoU on nuScenes [6]. Furthermore, it achieves the top score on the recent ScanNet++ benchmark [69], outperforming the next best approach by $\texttt{+}3.0$ mIoU. We call this approach DINO In The Room (DITR).
@@
Moreover, we show that even if images are not available during inference, it is still possible to utilize features from the same frozen DINOv2 model to enhance segmentation performance by pretraining 3D models through distillation [17]. Specifically, we take 3D point cloud datasets that have corresponding images and teach a 3D student model to output features aligned with those extracted from a DINOv2 teacher model, using a distillation objective (see Fig. 1 (b)). Subsequently, the pretrained, distilled student model can simply be fine-tuned for 3D segmentation in a regular fashion, without requiring corresponding images, and thus without causing any overhead during inference. This distillation pretraining enables the 3D model to capture the semantic richness of 2D foundation models without requiring any labeled data, effectively using them as distillation targets instead of semantic labels. Furthermore, it allows pretraining across multiple datasets without adjusting for dataset-specific semantic label sets, as the target feature space remains consistent. With this setup, we observe consistent improvements on all datasets compared to random initialization. On SemanticKITTI [5], where there is only a single image per 3D LiDAR scan, and on S3DIS [1], where camera views are sparser, this distillation setup even outperforms DITR. We call this approach Distill DITR (D-DITR).
@@
Altogether, these results highlight the significant yet underexplored potential of VFMs for 3D segmentation tasks. Therefore, we recommend that whenever corresponding images are available, they should be used to complement 3D segmentation models, preferably using injection and otherwise using distillation. Additionally, we note that while we focus on DINOv2 in this paper, as it is one of the strongest existing VFMs, we show that the general setup can also leverage even stronger models, like the very recently introduced DINOv3 [52].
@@
  • We demonstrate that DINOv2 can also act as a teacher to pretrain 3D models through distillation, improving performance without requiring 2D data during inference.
@@ 2 Related Work
3D Semantic Segmentation.  Despite architectural differences, most 3D backbones for dense tasks such as segmentation follow a U-Net-like hierarchical encoder–decoder design with skip connections. Early methods apply point-wise fully-connected layers [41, 42], or continuous 3D convolutions [38, 58] to capture local geometric patterns. Later, MinkUNet [9] significantly improves efficiency and accuracy by voxelizing point clouds and utilizing sparse 3D convolutions [15]. Following the success of Vision Transformers [12], attention mechanisms have been increasingly adopted for 3D segmentation, although their quadratic computational complexity poses challenges for large-scale scenes. To address this, several methods [30, 60, 68, 73, 63] restrict attention to local neighborhoods, where points are subsampled and aggregated hierarchically across multiple layers. Following this, Point Transformer V3 (PTv3) [62] employs space-filling curves to map 3D point clouds into 1D sequences, effectively preserving the local neighborhood structure while significantly simplifying and accelerating local attention computations. Most recently, Sonata [61] scales up PTv3, tripling the number of parameters, and adopts DINOv2-style self-supervised learning on point clouds, enabling the use of a substantially larger 3D training corpus [3, 74, 45, 4, 1, 10, 69], albeit still orders of magnitude smaller than its 2D counterparts. In contrast, our approach enhances PTv3 by directly incorporating features from powerful 2D VFMs such as DINOv2, thereby leveraging the vast amount of 2D training data on which these models are pretrained.
@@
While these works effectively use 2D information for 3D segmentation and achieve modest improvements, we observe that the community has not leveraged 2D VFMs for 3D segmentation, leaving a powerful information source untapped. In this work, to the best of our knowledge, we are the first to explore how the power of VFMs can be leveraged for state-of-the-art 3D segmentation, and demonstrate how 2D-to-3D injection with DITR yields new state-of-the-art results. Moreover, DITR is the first fusion approach shown to be effective for both indoor and outdoor scenes, highlighting its general applicability.
2D–3D Distillation.  Recent methods explore various strategies to distill 2D representations into 3D backbones for enhanced representation learning in LiDAR point clouds. Yan et al. [66] propose a joint training framework, where a 3D model is trained for both 3D segmentation and alignment with a 2D model. Several approaches apply contrastive learning either on super-points and super-pixels generated by pretrained 2D backbones [49, 32], or on prototypes created in both the 2D and 3D domains [8]. Others [35, 34] propose new contrastive loss functions, and conduct experiments with linear probing or limited-data setups. Recently, ScaLR [40] studies the effects of the size of the backbones and datasets in an image-to-LiDAR distillation setting. These methods focus on—and succeed in—image-to-LiDAR distillation for zero-shot, limited-data, or linear probing 3D segmentation settings. However, they often yield little or no improvement when used as pretrained models that are subsequently fine-tuned on 100 % of the data, and fall short of state-of-the-art methods in that setting. In contrast, we demonstrate that our D-DITR distillation is an effective pretraining step that significantly improves the fine-tuning performance of a state-of-the-art model (PTv3). Moreover, unlike previous methods, it also works especially well for dense indoor scenes.
@@ 3 Method
Figure 2: DITR architecture overview. We extract 2D image features from a frozen DINOv2 [36] model  and unproject them (2D-to-3D) onto the 3D point cloud. The unprojected features are subsequently max-pooled to create a multi-scale feature hierarchy. The raw point cloud is fed through a 3D backbone  and the unprojected image features are added to the skip connection between the encoder $\mathcal{E}_{l}$ and decoder $\mathcal{D}_{l}$ blocks on each level $l\in\{1,2,\dots,L\}$. The model is then trained with the regular segmentation loss.
@@ 3.1 Injection
In many modern 3D datasets, images from calibrated cameras are provided alongside the point clouds [10, 1, 6, 5, 55]. We leverage these images to inject semantically rich 2D features into the 3D backbone. An overview of this process is shown in Fig. 2: we first map 3D points to their corresponding pixels to extract associated DINOv2 features (2D-to-3D Mapping), and then inject these features into the skip connections of the 3D backbone’s decoder (3D Feature Fusion).
2D-to-3D Mapping.  Let $\mathcal{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ be a 3D point cloud of $N$ points, and assume a collection of $K$ calibrated cameras, $\{(\mathbf{I}_{k},\mathbf{K}_{k},\mathbf{T}_{k})\}^{K}_{k=1}$. Each camera has intrinsics $\mathbf{K}_{k}$ and world-to-camera extrinsics $\mathbf{T}_{k}$, with $\mathbf{I}_{k}$ denoting its captured image of resolution $H\times W$. We feed each 2D image $\mathbf{I}_{k}$ into a frozen DINOv2 ViT [12], obtaining patch-level embeddings $\mathbf{F}_{k}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D_{\text{2D}}}$, where $P$ is the patch size and $D_{\text{2D}}$ is the feature dimension. For each point $\mathbf{p}_{i}$, we transform it into the $k$-th camera space as $\mathbf{q}_{i}^{k}=\mathbf{T}_{k}\,(\mathbf{p}_{i},1)$ and multiply by the intrinsics $\mathbf{K}_{k}$ to obtain homogeneous pixel coordinates $\left(x_{i}^{k},\,y_{i}^{k},\,z_{i}^{k}\right)=\mathbf{K}_{k}\mathbf{q}_{i}^{k}$. The pixel coordinates are then given by $\left(u_{i}^{k},\,v_{i}^{k}\right)=\left(x_{i}^{k}/z_{i}^{k},\,y_{i}^{k}/z_{i}^{k}\right)$. We consider point $\mathbf{p}_{i}$ visible in the $k$-th image if $u_{i}^{k}\in[0,W)$, $v_{i}^{k}\in[0,H)$ and the depth $z_{i}^{k}$ is positive (view-frustum culling). If $\mathbf{p}_{i}$ is visible in the $k$-th image, we determine the corresponding patch index as
@@
3D Feature Fusion.  Because point clouds are unstructured, common 3D backbones [62, 9, 75] voxelize the input, keeping only one point per voxel, thereby reducing $\mathcal{P}$ to $\mathcal{P}^{\prime}$ with $M\leq N$ points. These backbones typically follow a U-Net-like encoder–decoder architecture, processing $\mathcal{P}^{\prime}$ at multiple spatial resolution levels $l\in\{1,2,\dots,L\}$. At each encoder level, a pooling layer downsamples the point features to a coarser spatial resolution. Formally, let $\mathbf{X}_{l}^{\mathcal{E}}\in\mathbb{R}^{M_{l}\times D_{l}}$ denote the output of the $l$-th level encoder block $\mathcal{E}_{l}$, where $l=1$ is the finest (original) resolution with $M_{1}=M$. After the encoder and the bottleneck $\mathcal{B}$, the decoder upsamples and refines these features back toward the original resolution, yielding $\mathbf{X}_{l}^{\mathcal{D}}$ as the output of each decoder block $\mathcal{D}_{l}$. To simplify notation, we define $\mathbf{X}_{L+1}^{\mathcal{D}}$ to be the output of the bottleneck $\mathcal{B}$.
@@
where ++ denotes element-wise addition, and each $f_{l}^{\cdot}$ is a linear projection into the decoder’s feature dimension followed by batch normalization [21] and GELU [16]. Finally, the output $\mathbf{X}^{\mathcal{D}}_{1}$ of the last decoder block yields per-point features for $\mathcal{P}^{\prime}$, which are passed through a linear segmentation head to produce class logits. With our injection approach, we ensure that rich 2D VFM features are available at all 3D decoder feature resolutions, enabling access to 2D information where needed and at different levels of granularity.